In our population of men, height is normally distributed with 70 inches(5'10') being the mean height. What is the probability that a randomly selected man will be 74 inches(6'2') or taller.
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

In our population of men, height is normally distributed with 70 inches(5'10') being the mean height. What is the probability that a randomly selected man will be 74 inches(6'2') or taller.
----------------
z(74) = (74-70)/2 = 2
---
P(x > 74) = P(z > 2) = normalcdf(2,100) = 0.0228
----
==============
Cheers,
Stan H.
==============

Answer by solver91311(24713)



No way to tell. You didn't provide the standard deviation value.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism